The mesa Redbirds football team plays in a stadium with seating capacity of 80,000. However, during the past season, attendance averaged only 50,000. The average ticket price was $30. If the price elasticity is -4, what would the team have to charge in order to fill the stadium?If the price were to be decreased to $27 and the average attendance increased to 60,000, what is the price elasticity?



1
Expert's answer
2021-10-02T10:49:02-0400

"P_1 = 30 \\
Q_1 = 50000 \\
Q_2 = 80000 \\
-4.0 = E_d = \frac{\frac{ΔQ}{\frac{Q_1+Q_2}{2}}}{\frac{ΔP}{\frac{P_1+P_2}{2}}} \\
= \frac{\frac{80000-50000}{50000-80000}}{\frac{P_2-30}{30+P_2}} \\
= \frac{30000}{13000} \times \frac{30+P_2}{P_2-30} \\
= 0.23076 \times \frac{30+P_2}{P_2-30} \\
-17.334 = \frac{30+P_2}{P_2-30} \\
-17.334P_2 + 520 = 30 +P_2 \\
490 = 18.334P_2 \\
P_2 = 26.726"